OpenStudy (bossimbacon):

What is the value of the expression? 2x3 + 5x2 will give fan and medal

OpenStudy (bossimbacon):

@Compassionate

OpenStudy (dangerousjesse):

I can't tell if those are x's or \(\times\)'s

OpenStudy (dangerousjesse):

If they're variables, it's \(7x^5\)

OpenStudy (dangerousjesse):

If they're symbols, it's \(2\times 3=x\) and \(5\times 2=y\) so \(x+y=\textsf{the answer}\)

OpenStudy (anonymous):

answer 16 yah welcs medal please

OpenStudy (anonymous):

@Bossimbacon

OpenStudy (dangerousjesse):

@kfullwood1 Don't give out answers.

OpenStudy (bossimbacon):

@kfullwood1 the 2x3 the 2 is an exponent same for 5x2

OpenStudy (bossimbacon):

the 2 is an exponent

OpenStudy (bossimbacon):

i mean the 3 is the exponent @kfullwood1